OpenStudy (loser66):

What is a complex polynomial? What is a even/odd complex polynomial? Please, help

OpenStudy (anonymous):

Polynomial with complex coefficients, I think.

OpenStudy (loser66):

That's it? I search but have no result. :(

OpenStudy (anonymous):

Even means \(f(-x) = f(x)\) and odd means \(f(-x) = -f(x)\)

OpenStudy (anonymous):

In either case, it must be \(f(0) = 0\) for a function to be even or odd.

OpenStudy (loser66):

My problem is : Prove either even or odd complex polynomial is a subspace of complex polynomial vector space. :( But I even NOT know what they are)

OpenStudy (loser66):

@wio, give me some suggestion, please

OpenStudy (anonymous):

I just gave you the definitions of what they are.

OpenStudy (anonymous):

Now you need to use the definition of vector subspace

OpenStudy (loser66):

yes, thanks for that, I will work on the proving problem by + and *

OpenStudy (loser66):

I don't see the different between real and complex coefficients in polynomials in the case. even/odd polynomial relate to variable only, right? f(-x) = -f(x)

OpenStudy (loser66):

and the coefficient can change the value of the function, not change the property of odd/even odd + odd = odd in both real and complex polynomial, right? so, why do they ask me to prove it in complex ? is there any thing different? I doubt myself.

OpenStudy (loser66):

scalar k*even = even, that's it. That's why I make a dumb question like this.

OpenStudy (anonymous):

I dunno

OpenStudy (anonymous):

Likely the proof for the real case is very similar.

OpenStudy (loser66):

yeah!!! I confused everything:(

OpenStudy (loser66):

Thanks for being here @wio